OpenStudy (anonymous):

Evaluate the expression when y = 4. 5 + y + 8

OpenStudy (welshfella):

replace y by 4:- 5 + 4 + 8 = ?

OpenStudy (anonymous):

A. 72 B. 21 C. 17 these are the answers

OpenStudy (welshfella):

I'm sure you can add 5 + 4 + 8 what is 5 + 4 equal to?

OpenStudy (anonymous):

ok i got it

OpenStudy (welshfella):

good

OpenStudy (nathalie0918):

Its C.17 5+4=9 9+8=17